1.Write the equation of the line with a slope of -1/3 and passing through the point (6, -4).
2.Write the equation of the line passing through the points (0, -4) and (-2, 2).
3.Write the equation of the line passing through the points (-6,1) and (-4,2).
4.Write the equation of the line with an undefined slope, passing through the point (2, 5).

Found 2 solutions by josgarithmetic, stanbon:
Answer by josgarithmetic(39838)

1:
Use those directly in point-slope form.

2 & 3:
Formula for slope of two points, and then choose either point, and use the point-slope equation form.

4:
Undefined slope means vertical line, or no change in x. x is a constant, and y has no restriction.

Answer by stanbon(75887)

1.Write the equation of the line with a slope of -1/3 and passing through the point (6, -4).
Form:: y = mx+b
-4 = (-1/3)6 + b
b = -2
y = (-1/3)x -2
-------------------------------
2.Write the equation of the line passing through the points (0, -4) and (-2, 2).
slope = (2--4)/(-2-0) = -3
y = mx + b
y = -3x - 4
---------------------
3.Write the equation of the line passing through the points (-6,1) and (-4,2).
m = (2-1)/(-4-6) = -1/10
y = mx + b
1 = (-1/10)(-6) + b
b = 4/10
y = (-1/10)x + (4/10)
----------------------
4.Write the equation of the line with an undefined slope, passing through the point (2, 5).
Undefined means the line is vertical
Ans: x = 2
-------------------
Cheers,
Stan H.
--------------
------